# SAGE: Spatial-visual Adaptive Graph Exploration for Visual Place Recognition

### v2

@@ Abstract
Visual Place Recognition (VPR) requires robust retrieval of geotagged images despite large appearance, viewpoint, and environmental variation. Prior methods focus on descriptor fine-tuning or fixed sampling strategies yet neglect the dynamic interplay between spatial context and visual similarity during training. We present SAGE (Spatial-visual Adaptive Graph Exploration), a unified training pipeline that enhances granular spatial–visual discrimination by jointly improving local feature aggregation, organize samples during training, and hard sample mining. We introduce a lightweight Soft Probing module that learns residual weights from training data for patch descriptors before bilinear aggregation, boosting distinctive local cues. During training we reconstruct an online geo–visual graph that fuses geographic proximity and current visual similarity so that candidate neighborhoods reflect the evolving embedding landscape. To concentrate learning on the most informative place neighborhoods, we seed clusters from high-affinity anchors and iteratively expand them with a greedy weighted clique expansion sampler. Implemented with a frozen DINOv2 backbone and parameter-efficient fine-tuning, SAGE achieves SOTA across eight benchmarks. It attains 98.9%, 95.8%, 94.5%, and 96.0% Recall@1 on SPED, Pitts30k-test, MSLS-val, and Nordland, respectively. Notably, our method obtains 100% Recall@10 on SPED only using 4096D global descriptors. Code and models will be released upon acceptance.
@@ 1 Introduction
Visual Place Recognition (VPR) matches a query image to its corresponding location within a large-scale geotagged database, serving as a fundamental capability for critical applications such as autonomous robot navigation (li2024navigation), loop closure detection for autonomous driving (Bench2drive), and large-scale map construction (ho2024map). The main challenge of VPR is maintaining robust retrieval performance under severe and unconstrained environmental changes, including extreme viewpoint shifts, illumination variations, adverse weather, long-term temporal drift, and frequent dynamic occluders, among others (tzachor2024effovpr; wang2025feature).
Early VPR methods relied on hand-crafted local descriptors (SIFT; SURF) and aggregated them into global encodings via pooling schemes such as Bag of Words or VLAD (BoW; vlad). However, these methods lack adaptability and perform poorly under large-scale appearance changes. With the advent of deep learning, learnable aggregation modules (netvlad; gem) were introduced, which enhanced descriptor compactness and robustness by learning task-specific pooling strategies. Subsequent research has mainly focused on simplifying, regularizing, or refining aggregation mechanisms to improve generalization and computational efficiency (jin2025edtformer), such as reducing reliance on explicit cluster centers or alleviating the “burstiness” problem of local features (supervlad; VLAD-BuFF).
@@
Recently, the advent of Visual Foundation Models (VFMs) (vit; dinov2) has advanced VPR by enabling the capture of long-range semantic dependencies and richer interactions between image patches (boq; FoL). These strategies reduce sensitivity to occlusion and background clutter and improve robustness with controlled parameter overhead. While recent adaptation methods for VFMs are notably efficient, different strategies within the broader VFMs ecosystem vary in resource consumption (EDTformer). For instance, fine-tuning a backbone’s encoder layers typically requires more computational resources than Parameter-Efficient Fine-Tuning (PEFT) approaches (jia2022visual; emvp). Furthermore, While the model’s understanding is dynamic, a pre-defined or static sampling policy (salad-cm) may fail to consistently present the most informative examples as learning progresses.
Recent research has explored constructing training batches that reflect real-world difficulty. However, a common strategy operates on a static “think-once, act-always” principle, relying on offline computations like pre-defined clustering based on initial features (QAConv-GS; GCL; salad-cm). This approach overlooks a crucial fact: what constitutes a hard sample is not a fixed property but a dynamic state that evolves together with the model’s embedding geometry during training. Effective mining thus requires an architectural “slow thinking” involving an iterative reassessment of difficulty. Without this, static methods quickly become obsolete. They continue feeding the model stale examples as old challenges turn trivial and new ones emerge at the decision boundary. This mismatch between a static sampling strategy and a dynamic learning process creates a critical bottleneck, hindering the model’s full discriminative potential.
@@
  • Efficient SOTA Accuracy. Implemented with a frozen DINOv2 backbone and parameter-efficient fine-tuning, SAGE sets SOTA on eight VPR benchmarks (Fig. 1), retaining competitiveness even with compact descriptors.
@@ 2 Related Work
Visual Place Recognition (VPR) requires global descriptors to remain compact and robust under substantial variations in viewpoint, illumination, and scene structure. Early methods for generating global descriptors, such as NetVLAD (netvlad), utilized a vast set of cluster centers, which rendered them vulnerable to domain shifts. Although self-supervised hard sample mining (sfrs) has been introduced to mitigate this issue, such methods remain constrained by the representational capacity of CNN backbones, yielding suboptimal performance. With the incorporation of spatial attention (delf) and feature reweighting (solar), descriptor robustness has been yet improved. Other works focus on optimizing feature aggregation (mixvpr; salad; boq), which improves performance but relies on fixed aggregation strategies and thus lacks adaptability to dynamically evolving embeddings. Recent studies have highlighted the effectiveness of modulating feature magnitudes in a lightweight, data-dependent manner prior to their aggregation. Such approaches include non-local attention for adaptive spatial weighting (chen2023self), efficient context encoding (huang2022efficient), parameter-efficient tuning with second-order moments (gao2023tuning), and the exponentially weighted fusion of pooling kernels (stergiou2022adapool). Other methods reduce overhead by lowering the number of clusters or entirely eliminating clustering through centroid-free probes (CFP) (supervlad; emvp), yielding compact descriptors by utilizing second-order feature statistics. Two-stage approaches (selavpr; FoL) improve retrieval accuracy through local feature re-ranking but introduce additional computational overhead. Recent single-stage methods can achieve comparable or even superior performance using only global features (megaloc; superplace). This pursuit of efficiency is also reflected in adapting powerful Visual Foundation Models (VFMs) for downstream tasks (zhang2025d; embodiedplace). With the proliferation of VFMs, Parameter-Efficient Fine-Tuning (PEFT) has emerged as a crucial paradigm. Instead of fine-tuning the entire backbone, these methods update only a small subset of parameters, such as lightweight adapters or normalization layers, thereby significantly enhancing training efficiency (jia2022visual; emvp). Our method follows this paradigm by freezing the backbone while introducing lightweight modules to enhance feature discriminability. Cross-image correlation methods (cricavpr; emvp) further enhance matching by capturing and modeling inter-image dependencies. In line with this paradigm, our approach strategically enhances feature discriminability. We introduce Soft Probe, a lightweight residual module that adaptively amplifies salient local regions, and InteractHead, which models cross-image dependencies. Together, they significantly boost descriptor quality with minimal parameter overhead.
@@
In deep metric learning, dynamic sampling strategies (DynamicSampling) adjust the importance of training pairs with epoch-dependent weighting terms, organizing them in an “easy-to-hard” order, which enables the network to first learn general category boundaries from easy samples and then focus on hard samples in later stages. More recently, a graph-based sampling method (QAConv-GS) has been proposed, which constructs a nearest neighbor graph from class embeddings at the beginning of each epoch. By selecting an anchor class and its neighboring classes to form training batches, this approach improves the discriminative power of learned embeddings and enhances training efficiency. In VPR, spatial graphs have been used to encode geographic relationships. For example, the MMS-VPR benchmark (MMSVPR) represents street intersections and road segments as nodes and edges, leveraging topological context to improve retrieval performance. Such approaches often rely on mining discriminative regions or focusing on hard positive samples (aanet; seidenschwarz2021learning; fang2022adversarial; FoL) to enhance accuracy and robustness, but ignore geo-information. Other works reformulate VPR as a classification task (cosplace; eigenplaces) to avoid explicit mining, but these remain limited by static sample selection and the neglect of geographic information. Moreover, on sparse datasets such as GSV-Cities (gsv), these limitations restrict generalization capability. On dense datasets such as MSLS (msls), offline clustering methods (salad-cm) partition visually similar and geographically neighboring images into fixed clusters for training. Another category of approaches leverages hybrid strategies (kalantidis2020hard), hard negative mining (garg2022seqmatchnet; deuser2023sample4geo; ali2023global) or generation (peng2024globally) to enhance retrieval performance. However, most of these approaches struggle to generalize effectively. In contrast to such static or scheduled strategies, SAGE reconstructs a geo-visual graph at every training epoch and employs greedy sampling to focus on the most densely populated and challenging clusters in the evolving embedding spaces, achieving superior performance across datasets.
@@ 3.1 Feature Extraction
DINOv2 provides strong visual representations from large-scale self-supervised pretraining. We use a pretrained DINOv2 as a frozen backbone and achieve parameter-efficient fine-tuning by inserting learnable Dynamic Power Normalization (DPN) layers into the last $N$ encoder blocks (dinov2; emvp). As shown in Fig. 2(a), an input image $I\in\mathbb{R}^{H\times W\times 3}$ passes through two stages. Uncalibrated Blocks extract base features, and Recalibrated Blocks integrate DPN to recalibrate features and produce task-specific representations. The backbone outputs one learnable class token and $L$ patch tokens, forming a token matrix $\mathbf{f}\in\mathbb{R}^{(L+1)\times M}$, where $M$ is the embedding dimension. SoftP and InteractHead then aggregate and refine these tokens to strengthen cross-image correspondence and yield a discriminative global descriptor $\mathbf{F}\in\mathbb{R}^{D\times K}$ for place recognition.
@@ 3.2 Soft Probing
Centroid-Free Probing (CFP) crafts a global image descriptor $\mathbf{f}$ by aggregating the $L$ local spatial descriptors $\{X_{i}\}_{i=1}^{L}$ into a second-order Gram-like covariance matrix, which eliminates the need for offline cluster centers (emvp; supervlad). Although related methods like Moment Probing also leverage second-order statistics (wang2022dropcov; gao2023tuning), a key limitation of CFP is its uniform treatment of all descriptors. This equal weighting can underemphasize subtle yet discriminative local cues. To address this shortcoming, we introduce Soft Probing (SoftP), a lightweight module shown in Fig. 2(d), which adaptively emphasizes informative spatial locations while preserving the underlying feature geometry before the final aggregation step.
@@
This residual reweighting behaves like a soft, data-driven attention mechanism: it amplifies salient responses while avoiding destructive rescaling of channel structure (solar). Under mild assumptions (or to first order when $\beta_{i}$ is small and the mean shift is negligible) (gao2023tuning), the variance of each dimension of the modulated descriptors increases approximately as:
@@ 3.3 Online Graph Creation
To prepare for the graph creation, the image descriptors $\mathbf{f}_{i}\in\mathbb{R}^{D\times K}$ produced by SoftP are first processed by the InteractHead. Departing from prior method that partition features via learned cluster assignments(cricavpr; supervlad), InteractHead deterministically splits each descriptor into $S$ fixed length segments, $\{\mathbf{f}_{i}^{(s)}\}_{s=1}^{S}$. To enable cross-image attention, these segments are rearranged such that for each index $s$, the segments from all $B$ images form a sequence. These sequences are then processed by a two-layer Transformer encoder ($\mathcal{E}$) with GELU activation. This encoder structure applies attention across the batch for each segment type, capturing consistent correlations across views and improving descriptor robustness. The enhanced descriptors $\mathbf{F}$ are obtained by:
@@ 4.1 Datasets and Performance Evaluation
We validate SAGE on a diverse collection of VPR benchmarks (Tab. 1) covering common real-world challenges: Pitts30k-test and Pitts250k-test (large viewpoint variation) (pitts), SPED (low-quality / high scene depth and condition changes) (SPED), MSLS-val (multi-year urban/suburban variability) (msls), Nordland (four-season natural scenes) (nordland), Tokyo247 (multi-view urban captures) (tokyo247), AmsterTime (historical grayscale vs. contemporary RGB) (amstertime), and Eynsham (rural grayscale route) (eynsham; berton2022deep). Further details can be found in App. A.3.
The experiment adopts Recall@N (R@N) as the evaluation metric, i.e., the percentage of query images for which at least one of the top-N retrieved database images geographically matches the query image (within a preset threshold). Thresholds follow standard protocols: 25 meters for Pitts30k-test, Pitts250k-test, Tokyo24/7, Eynsham and SPED; MSLS-val uses 25 meters with azimuth within 40 degrees; Nordland uses $\pm 10$ frames (pitts; SPED; msls).

### v3

# SAGE: Spatial-visual Adaptive Graph Exploration for Efficient Visual Place Recognition

@@ Abstract
Visual Place Recognition (VPR) requires robust retrieval of geotagged images despite large appearance, viewpoint, and environmental variation. Prior methods focus on descriptor fine-tuning or fixed sampling strategies yet neglect the dynamic interplay between spatial context and visual similarity during training. We present SAGE (Spatial-visual Adaptive Graph Exploration), a unified training pipeline that enhances granular spatial–visual discrimination by jointly improving local feature aggregation, organize samples during training, and hard sample mining. We introduce a lightweight Soft Probing module that learns residual weights from training data for patch descriptors before bilinear aggregation, boosting distinctive local cues. During training we reconstruct an online geo–visual graph that fuses geographic proximity and current visual similarity so that candidate neighborhoods reflect the evolving embedding landscape. To concentrate learning on the most informative place neighborhoods, we seed clusters from high-affinity anchors and iteratively expand them with a greedy weighted clique expansion sampler. Implemented with a frozen DINOv2 backbone and parameter-efficient fine-tuning, SAGE achieves SOTA across eight benchmarks. Notably, our method obtains 100% Recall@10 on SPED only using 4096D global descriptors. The code and model are available at https://github.com/chenshunpeng/SAGE.
@@ 1 Introduction
Visual Place Recognition (VPR) matches a query image to its corresponding location within a large-scale geotagged database, serving as a fundamental capability for critical applications such as autonomous robot navigation (Han et al., 2025), loop closure detection for autonomous driving (Teng et al., 2026), and large-scale map construction (Zhu et al., 2024). The main challenge of VPR is maintaining robust retrieval performance under severe and unconstrained environmental changes, including extreme viewpoint shifts, illumination variations, adverse weather, long-term temporal drift, and frequent dynamic occluders, among others (Liu et al., 2024; Zhu et al., 2025).
Early VPR methods relied on hand-crafted local descriptors (Lowe, 2004; Bay et al., 2008) and aggregated them into global encodings via pooling schemes such as Bag of Words or VLAD (Angeli et al., 2008; Jégou et al., 2010). However, these methods lack adaptability and perform poorly under large-scale appearance changes. With the advent of deep learning, learnable aggregation modules (Arandjelovic et al., 2016; Radenović et al., 2018) were introduced, which enhanced descriptor compactness and robustness by learning task-specific pooling strategies. Subsequent research has mainly focused on simplifying, regularizing, or refining aggregation mechanisms to improve generalization and computational efficiency (Jin et al., 2025b), such as reducing reliance on explicit cluster centers or alleviating the “burstiness” problem of local features (Lu et al., 2024d; Khaliq et al., 2024).
@@
Recently, the advent of Visual Foundation Models (VFMs) (Dosovitskiy et al., 2020; Oquab et al., 2023) has advanced VPR by enabling the capture of long-range semantic dependencies and richer interactions between image patches (Ali-bey et al., 2024; Lu et al., 2025). These strategies reduce sensitivity to occlusion and background clutter and improve robustness with controlled parameter overhead. While recent adaptation methods for VFMs are notably efficient, different strategies within the broader VFMs ecosystem vary in resource consumption (Jin et al., 2025a). For instance, fine-tuning a backbone’s encoder layers typically requires more computational resources than Parameter-Efficient Fine-Tuning (PEFT) approaches (Jia et al., 2022; Qiu et al., 2024). Furthermore, while the model’s understanding is dynamic, a pre-defined or static sampling policy (Izquierdo and Civera, 2025) may fail to consistently present the most informative examples as learning progresses.
Recent research has explored constructing training batches that reflect real-world difficulty. However, a common strategy operates on a static “think-once, act-always” principle, relying on offline computations like pre-defined clustering based on initial features (Liao and Shao, 2022; Leyva-Vallina et al., 2023; Tzachor et al., 2025; Wang et al., 2025; Izquierdo and Civera, 2025). Such approaches often treat geographic priors and visual similarity independently, overlooking a crucial fact: what constitutes a real-world hard sample is governed by the intricate, dynamic interplay between spatial proximity and visual ambiguity. This hardness is not a fixed property, but a state that evolves with the model’s embedding geometry during training. Effective mining thus requires an architectural “slow thinking” involving an iterative reassessment of difficulty. Without this, static methods gradually become obsolete. They continue feeding the model stale examples, wasting computational resources and limiting learning efficiency, as old challenges turn trivial and new ones emerge at the decision boundary. Ultimately, the inability to concurrently enhance granular local expressiveness and dynamically align the sampling process with the evolving geo-visual landscape leaves a significant gap in current training paradigms, potentially limiting their adaptability in real-world scenarios.
@@
  • Efficient SOTA Accuracy. Implemented with a frozen DINOv2 backbone and parameter-efficient fine-tuning, SAGE sets a new SOTA across eight challenging VPR benchmarks (Fig. 1), retaining robust retrieval competitiveness even with highly compact descriptors for large-scale deployment.
@@ 2 Related Work
Visual Place Recognition (VPR) requires global descriptors to remain compact and robust under substantial variations in viewpoint, illumination, and scene structure. Early methods for generating global descriptors, such as NetVLAD (Arandjelovic et al., 2016), utilized a vast set of cluster centers, which rendered them vulnerable to domain shifts. Although self-supervised hard sample mining (Ge et al., 2020) has been introduced to mitigate this issue, such methods remain constrained by the representational capacity of CNN backbones, yielding suboptimal performance. With the incorporation of spatial attention (Noh et al., 2017) and feature reweighting (Ng et al., 2020), descriptor robustness has been yet improved. Other works focus on optimizing feature aggregation (Ali-Bey et al., 2023b; Izquierdo and Civera, 2024; Ali-bey et al., 2024), which improves performance but relies on fixed aggregation strategies and thus lacks adaptability to dynamically evolving embeddings. Recent studies have highlighted the effectiveness of modulating feature magnitudes in a lightweight, data-dependent manner prior to their aggregation. Such approaches include non-local attention for adaptive spatial weighting (Chen et al., 2023), efficient context encoding (Huang et al., 2022), parameter-efficient tuning with second-order moments (Gao et al., 2023), and the exponentially weighted fusion of pooling kernels (Stergiou and Poppe, 2022). Other methods reduce overhead by lowering the number of clusters or entirely eliminating clustering through centroid-free probes (CFP) (Lu et al., 2024d; Qiu et al., 2024), yielding compact descriptors by utilizing second-order feature statistics. Two-stage approaches (Wang et al., 2025; Lu et al., 2024a) improve retrieval accuracy through local feature re-ranking but introduce additional computational overhead. Recent single-stage methods can achieve comparable or even superior performance using only global features (Berton and Masone, 2025; Liu et al., 2025b). This pursuit of efficiency is also reflected in adapting powerful Visual Foundation Models (VFMs) for downstream tasks (Zhang et al., 2025; Liu et al., 2025a). With the proliferation of VFMs, Parameter-Efficient Fine-Tuning (PEFT) has emerged as a crucial paradigm. Instead of fine-tuning the entire backbone, these methods update only a small subset of parameters, such as lightweight adapters or normalization layers, thereby significantly enhancing training efficiency (Jia et al., 2022; Qiu et al., 2024). Our method follows this paradigm by freezing the backbone while introducing lightweight modules to enhance feature discriminability. Cross-image correlation methods (Lu et al., 2024b; Qiu et al., 2024) further enhance matching by capturing and modeling inter-image dependencies. In line with this paradigm, our approach strategically enhances feature discriminability. We introduce Soft Probe, a lightweight residual module that adaptively amplifies salient local regions, and InteractHead, which models cross-image dependencies. Together, they significantly boost descriptor quality with minimal parameter overhead.
@@
In deep metric learning, dynamic sampling strategies (Liang et al., 2021) adjust the importance of training pairs with epoch-dependent weighting terms, organizing them in an “easy-to-hard” order, which enables the network to first learn general category boundaries from easy samples and then focus on hard samples in later stages. More recently, a graph-based sampling method (Liao and Shao, 2022) has been proposed, which constructs a nearest neighbor graph from class embeddings at the beginning of each epoch. By selecting an anchor class and its neighboring classes to form training batches, this approach improves the discriminative power of learned embeddings and enhances training efficiency. In VPR, spatial graphs have been used to encode geographic relationships. For example, the MMS-VPR benchmark (Ou et al., 2025) represents street intersections and road segments as nodes and edges, leveraging topological context to improve retrieval performance. Such approaches often rely on mining discriminative regions or focusing on hard positive samples (Lu et al., 2023; Seidenschwarz et al., 2021; Fang et al., 2022; Wang et al., 2025) to enhance accuracy and robustness, but ignore geo-information. Other works reformulate VPR as a classification task (Berton et al., 2022a; 2023) to avoid explicit mining, but these remain limited by static sample selection and the neglect of geographic information. Moreover, on sparse datasets such as GSV-Cities (Ali-bey et al., 2022), these limitations restrict generalization capability. On dense datasets such as MSLS (Warburg et al., 2020), offline clustering methods (Izquierdo and Civera, 2025) partition visually similar and geographically neighboring images into fixed clusters for training. Another category of approaches leverages hybrid strategies (Kalantidis et al., 2020), hard negative mining (Garg et al., 2022; Deuser et al., 2023; Ali-Bey et al., 2023a) or generation (Peng et al., 2024) to enhance retrieval performance. However, most of these approaches struggle to generalize effectively. In contrast to such static or scheduled strategies, SAGE reconstructs a geo-visual graph at every training epoch and employs greedy sampling to focus on the most densely populated and challenging clusters in the evolving embedding spaces, achieving superior performance across datasets.
@@ 3.1 Feature Extraction
DINOv2 provides strong visual representations from large-scale self-supervised pretraining. We use a pretrained DINOv2 as a frozen backbone and achieve parameter-efficient fine-tuning by inserting learnable Dynamic Power Normalization (DPN) layers into the last $N$ encoder blocks (Oquab et al., 2023; Qiu et al., 2024). As shown in Fig. 2(a), an input image $I\in\mathbb{R}^{H\times W\times 3}$ passes through two stages. Uncalibrated Blocks extract base features, and Recalibrated Blocks integrate DPN to recalibrate features and produce task-specific representations. The backbone outputs one learnable class token and $L$ patch tokens, forming a token matrix $\mathbf{f}\in\mathbb{R}^{(L+1)\times M}$, where $M$ is the embedding dimension. SoftP and InteractHead then aggregate and refine these tokens to strengthen cross-image correspondence and yield a discriminative global descriptor $\mathbf{F}\in\mathbb{R}^{D\times K}$ for place recognition.
@@ 3.2 Soft Probing
Centroid-Free Probing (CFP) crafts a global image descriptor $\mathbf{f}$ by aggregating the $L$ local spatial descriptors $\{X_{i}\}_{i=1}^{L}$ into a second-order Gram-like covariance matrix, which eliminates the need for offline cluster centers (Qiu et al., 2024; Lu et al., 2024d). Although related methods like Moment Probing also leverage second-order statistics (Wang et al., 2022; Gao et al., 2023), a key limitation of CFP is its uniform treatment of all descriptors. This equal weighting can underemphasize subtle yet discriminative local cues. To address this shortcoming, we introduce Soft Probing (SoftP), a lightweight module shown in Fig. 2(d), which adaptively emphasizes informative spatial locations while preserving the underlying feature geometry before the final aggregation step.
@@
This residual reweighting behaves like a soft, data-driven attention mechanism: it amplifies salient responses while avoiding destructive rescaling of channel structure (Ng et al., 2020). Under mild assumptions (or to first order when $\beta_{i}$ is small and the mean shift is negligible) (Gao et al., 2023), the variance of each dimension of the modulated descriptors increases approximately as:
@@ 3.3 Online Graph Creation
To prepare for the graph creation, the image descriptors $\mathbf{f}_{i}\in\mathbb{R}^{D\times K}$ produced by SoftP are first processed by the InteractHead. Departing from prior method that partition features via learned cluster assignments(Lu et al., 2024b; d), InteractHead deterministically splits each descriptor into $S$ fixed length segments, $\{\mathbf{f}_{i}^{(s)}\}_{s=1}^{S}$. To enable cross-image attention, these segments are rearranged such that for each index $s$, the segments from all $B$ images form a sequence. These sequences are then processed by a two-layer Transformer encoder ($\mathcal{E}$) with GELU activation. This encoder structure applies attention across the batch for each segment type, capturing consistent correlations across views and improving descriptor robustness. The enhanced descriptors $\mathbf{F}$ are obtained by:
@@ 4.1 Datasets and Performance Evaluation
We validate SAGE on a diverse collection of VPR benchmarks (Tab. 1) covering common real-world challenges: Pitts30k-test and Pitts250k-test (large viewpoint variation) (Torii et al., 2013), SPED (low-quality / high scene depth and condition changes) (Chen et al., 2017), MSLS-val (multi-year urban/suburban variability) (Warburg et al., 2020), Nordland (four-season natural scenes) (Sünderhauf et al., 2013), Tokyo247 (multi-view urban captures) (Torii et al., 2015), AmsterTime (historical grayscale vs. contemporary RGB) (Yildiz et al., 2022), and Eynsham (rural grayscale route) (Cummins and Newman, 2010; Berton et al., 2022b). Further details can be found in App. A.3.
The experiment adopts Recall@N (R@N) as the evaluation metric, i.e., the percentage of query images for which at least one of the top-N retrieved database images geographically matches the query image (within a preset threshold). Thresholds follow standard protocols: 25 meters for Pitts30k-test, Pitts250k-test, Tokyo24/7, Eynsham and SPED; MSLS-val uses 25 meters with azimuth within 40 degrees; Nordland uses $\pm 10$ frames (Torii et al., 2013; Chen et al., 2017; Warburg et al., 2020).